Your parents have informed you that they plan to reward you for your academic achievements with $9000 when you graduate next year. The risk -free rate is 3%, and the expected return on the market is 11%, with a standard deviation of 35%. What is the present value of this gift if you assume that your future cash flow is independent of the return on the market? Round your answer to the nearest dollar. 8,738 9,270 8,108 9,000
Answer #1

future value = $9000

r=3%

n=1 year

present value=(future value)/(1+r)^n = 9000/(1+3%)^1 = $8738